# LAMMPS-KOKKOS: Performance Portable Molecular Dynamics Across Exascale Architectures

### v1

@@ Abstract.
Since its inception in 1995, LAMMPS has grown to be a world-class molecular dynamics code, with thousands of users, over one million lines of code, and multi-scale simulation capabilities. We discuss how LAMMPS has adapted to the modern heterogeneous computing landscape by integrating the Kokkos performance portability library into the existing C++ code. We investigate performance portability of simple pairwise, many-body reactive, and machine-learned force-field interatomic potentials. We present results on GPUs across different vendors and generations, and analyze performance trends, probing FLOPS throughput, memory bandwidths, cache capabilities, and thread-atomic operation performance. Finally, we demonstrate strong scaling on all current US exascale machines – OLCF Frontier, and ALCF Aurora, and NNSA El Capitan – for the three potentials.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Workshop on Performance, Portability & Productivity in HPC; November 16–21, 2025; St. Louis, MOisbn: 978-1-4503-XXXX-X/2018/06ccs: General and reference Performanceccs: Software and its engineering Ultra-large-scale systemsccs: Computing methodologies Parallel algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Distributed algorithmsccs: Computing methodologies Machine learningccs: Computing methodologies Molecular simulation
@@ 1. Introduction
Molecular dynamics is a method for studying materials and molecules by directly integrating Newton’s second law to obtain trajectories that describe chemical reactions, heat transport, diffusion, and other phenomena that occur at the time and length scales that are within computational reach of atomistic methods. Since the initial efforts by Rahman in 1964 (Rahman, 1964), the increase in applicability of molecular dynamics has been driven by the exponential growth in computational power that enables simulations of larger systems over longer timescales with higher accuracy (Plimpton and Thompson, 2012; Perez et al., 2025).
Further progress made in the 1990s and the spread of distributed computing presented a major leap. The natural spatial decomposition of molecular dynamics simulations made them a prime example for adopting the new massive parallelism paradigm, as demonstrated by the first releases of still-popular software such as LAMMPS in 1995 (Plimpton, 1995), as well as GROMACS (Abraham et al., 2015) and NAMD (Phillips et al., 2005). Over the next decade and a half, LAMMPS grew in scope and capability, from 23k lines of Fortran code in 1999, to 148k lines of C++ in 2009, but the MPI-distributed domain decomposition computing paradigm remained the same. Then, GPUs became more general purpose with the introduction of programming models such as CUDA, and their high FLOP rate and improved power efficiency gave significant advantages over CPUs, leading to widespread adoption of heterogeneous computing clusters.
With parallelism being a major strong point of GPUs, molecular dynamics codes with their large number of force calculations once again became an early adopter of the new heterogeneous computing paradigm. New codes were developed such as HOOMD-blue (Anderson et al., 2020) and HAL’s MD (Colberg and Höfling, 2011), while existing codes raced to adopt the new technology as quickly and efficiently as possible. For LAMMPS, whose modular package system (described in 3.1 below) allows development of different approaches simultaneously, two competing strategies soon emerged.
@@
This paper instead focuses on the second strategy: the KOKKOS package in LAMMPS, which uses the Kokkos library (Trott et al., 2021) to achieve performance portability. The KOKKOS package (released in 2014) has design origins in the USER-CUDA package developed by Trott (Trott, 2011; Trott et al., 2010), which was released as part of LAMMPS in 2011. The USER-CUDA package attempted to be “GPU resident”, meaning as many kernels as possible run on the GPU. This avoids copying data between GPU and CPU at every timestep, leading to performance benefits. However, data transfer is still necessary to maintain compatibility when using functionality not yet ported to GPUs. The KOKKOS package brought the additional promise of keeping LAMMPS vendor-agnostic, which has proven vital given that the current US exascale landscape is based on GPUs from AMD and Intel, and other large supercomputers contain GPUs from NVIDIA.
@@
In this paper, we discuss the current structure of LAMMPS and how the optional KOKKOS package provides seamless performance portability for the user when enabled. We include a comprehensive set of benchmarks for comparing how NVIDIA GPU performance has improved over hardware generations (V100-16GB-SXM3, A100-40GB-SXM4, and H100-HBM3-SXM5), and how LAMMPS benchmarks perform on the three existing exascale systems Frontier (AMD MI250X) (Top500, 2024d), El Capitan (AMD MI300A) (Top500, 2024b) and Aurora (Intel Data Center GPU Max 1550 - subsequently referred to as PVC) (Top500, 2024a), along with Eos (NVIDIA DGX H100 Superpod) (Top500, 2024c). We present case studies on the Lennard-Jones, SNAP and ReaxFF potentials, and a discussion of how LAMMPS keeps up with current developments in the field of molecular dynamics such as machine-learning-based potentials.
pair_style eampair_style eam/kkInput script:class PairEAMvirtual void compute(...)comm->reverse_comm(this);}int pack_reverse_comm(...)class PairEAMKokkosvoid compute(...){commKK->reverse_comm(this);}int pack_reverse_comm_kokkos(...)class Pairvirtual void compute(...)class KokkosBasevirtual int pack_reverse_comm_kokkos(...)class AtomVecAtomicdouble** xclass AtomVecAtomicKokkosDualView<double*[3]> xclass CommBrickvirtual void reverse_comm(Pair* pair){pair->pack_reverse_comm(...) } class CommKokkosvoid reverse_comm(Pair* pair){(KokkosBase*)pair-> pack_reverse_comm_kokkos(...) }
@@
Figure 1. Mapping (red arrows) from user input script commands to LAMMPS class hierarchy and how the classes interact within and outside of the KOKKOS package. The pair styles operate on atomic data stored as pointers in AtomVecAtomic and Kokkos::DualViews in AtomVecAtomicKokkos, with the pointer in the former aliased to the CPU mirror of the latter. The EAM pair style requires additional communication, which is performed with calls to the LAMMPS communication classes. Following Unified Modeling Language (ULM) conventions, solid black arrows indicate inheritance, while dashed arrows indicate usage.
@@
Table 1. GPU architecture properties, listing HBM memory bandwidth (“BW”) and capacity, floating point throughput (“FP64”, excludes matrix multiplication hardware), and cache size (single numbers are used for combined caches, two numbers for separate hardware and software managed caches). For AMD MI250 and Intel PVC a single logical GPU is listed, not the full package.
@@ 2. LAMMPS basic code structure
LAMMPS is one of the leading molecular dynamics simulation codes in part because of the flexibility and collection of capabilities that have grown through over 30 years of development and user contributions (Kohlmeyer and Berger, 2025; Thompson et al., 2022).
@@ 2.2. Styles
These persistent classes are referred to as styles, which belong to different categories described more closely in the LAMMPS reference paper (Thompson et al., 2022). LAMMPS contains hundreds of different styles, in different categories. “Fix” styles have methods that are called at arbitrary points and intervals during the simulation to either modify the trajectory of the simulation or generate output, while “compute” styles can be viewed as restricted version of fixes that only produce output accessible to the user within the input script, without modifying the state of the system. The case studies presented in 4 are examples of pair styles, with each pair style implementing a different model for the potential energy, often referred to as simply “potentials”. At each timestep, the derivatives of the selected potential with respect to atomic coordinates provide the forces that are used for time integration, making pair styles the most important category of styles since they are typically the most expensive part of a simulation. The examples in 4 span a range of complexities, from a simple pairwise potential to a state-of-the-art machine-learning-based potential.
3. LAMMPS KOKKOS package
The KOKKOS package in LAMMPS was one of the first serious applications for Kokkos, with some of Kokkos’s design features directly inspired by use-cases from LAMMPS – including in the area of data structures and parallel execution. In this section, we describe how Kokkos is integrated to seamlessly enable performance portability without losing access to pre-existing non-Kokkos functionality.
@@ 3.1. Packages and suffixes
The “accelerator” packages, on the other hand, provide replacement functionality in the form of accelerated styles, i.e., classes that provide equivalent functionality but improve computational performance on different hardware. This paper describes the KOKKOS package, which provides alternate versions of many classes that are accelerated by the Kokkos library for performance portability. As shown in 1, the KOKKOS package contains a PairEAMKokkos class for the EAM potential that is equivalent to the base PairEAM class from which it inherits, but its computational routines implement Kokkos library abstractions.
At runtime, the accelerated versions need to be selected through the map from input script command to C++ class described in 2.1 and 1. Each accelerated style is registered and added to the map using the same macro as the non-accelerated style, with the convention that the accelerated styles append a package-specific suffix, namely /kk for the KOKKOS package. The user can then selectively choose the accelerated version of a style using, for example, the eam/kk pair style instead of eam in their input script, or they can use the Kokkos-accelerated version of all possible styles used in the input script by specifying a global suffix. Through careful design of the data structures and data movement patterns described in this section, the user is allowed to accelerate the parts of their simulation for which accelerated styles are available, without losing access to those for which no Kokkos support currently exists.
@@ 3.2. Data Structures
The primary data structure in Kokkos is the Kokkos::View, a structure that can represent multi-dimensional arrays, and encodes both data layout and data accessible through type mechanism via its Layout and MemorySpace template parameters. Kokkos::View strongly informed the design of the ISO C++23 std::mdspan capability. The Kokkos variants of styles in LAMMPS generally contain host and device variants of data encapsulated in a Kokkos::DualView. This data structure contains a compatible Kokkos::View for both (i.e. CPU) and a device (i.e. GPU) memory space to help with unstructured data synchronization. In particular, it has functionality to keep track of when data was modified, and thus when data has to be synced. The classic non-Kokkos data fields in LAMMPS are initialized to alias the allocations underlying the host View of the Kokkos data structures via Kokkos’s View interoperability with raw pointers.
@@
Another higher level data structure relied upon by LAMMPS is ScatterView. This data structure was designed to handle unstructured accumulation of data from multiple threads in a way that write conflicts are avoided. It can transparently swap between using atomic operations, a data duplication strategy, or even simple sequential accumulation in case of non-threaded execution. This is important to handle the different amount of concurrency and atomic operation throughput on various architectures. On CPUs, data duplication with a subsequent combining step is often the most effective way to deal with write conflicts, while on GPUs data duplication is unfeasible due to the large number of active threads (O(100,000)) and thus atomic operations need to be used.
@@ 3.3. Execution Control
Nearly every Kokkos-based style in LAMMPS is templated on the device type and is instantiated for both the default host and device execution space simultaneously. This dual instantiation pattern enables users to request a Kokkos-based style for either host CPU or GPU device at runtime from the input script through the /kk/host suffix and the /kk/device suffix (equivalent to /kk, see 3.1), which can greatly improve execution control. One example of potential benefits is in the communication phase, where, depending on the size of the problem, the availability of GPU-aware MPI, and the actual hardware, it may be more performant to keep all communication routines (packing, unpacking, sending data) on host, or execute it on the device.
Kokkos parallel execution is controlled via execution patterns and execution policies. The primary patterns are parallel_for, parallel_reduce and parallel_scan which execute a callable for each index in an iteration space, and if applicable combine results via reductions. The main policies are RangePolicy, MDRangePolicy and TeamPolicy. RangePolicy and MDRangePolicy simply iterate over a 1D or multi-dimensional range of indices respectively. The latter also enables tiling, which can be beneficial to achieve better cache locality in multi-dimensional loop patterns.
TeamPolicy enables the use of hierarchical (nested) parallelism, i.e. exposing concurrency in non-tightly nested loops. Specifically it exposes a concept of thread teams, which can collaboratively execute nested parallel loops, as well as vector level parallelism as a third nesting level using TeamThreadRange, TeamVectorRange and ThreadVectorRange respectively. It also enables the use of scratch memory, which on GPUs can be mapped to software-managed caches such as NVIDIA’s “shared memory”, a topic which will be explored in more depth in 4.4.
@@ 4. Case studies
For a general pairwise potential, the total energy of the system is a sum over all pairs of atoms that are within a specified cutoff distance. The inter-atomic forces are then the pairwise sum of the derivative of the potential energy with respect to position. The summation over pairs of close atoms is enabled by the neighbor list constructed and periodically updated by LAMMPS according to the algorithms described in the LAMMPS reference paper (Thompson et al., 2022).
@@
(1) $E=\sum_{i<k,r_{ik}<r_{\text{c}}}\left(4\epsilon\left[\left(\sigma/r_{ik}\right)^{12}-\left(\sigma/r_{ik}\right)^{6}\right]\right)=\sum_{i<k,r_{ik}<r_{\text{c}}}U^{(2)}~(r_{ik}),$
with $r_{ik}$ the distance between atoms $i$ and $k$, $r_{\text{c}}$ is the interaction cutoff distance that may vary between atomic systems, and $\epsilon$ and $\sigma$ are coefficients unique to LJ that may also vary between given atomic systems. LJ refers to a specific functional form of the more general idea of a two body potential $U^{(2)}(r_{ik})$. Different systems call for different functional forms. LJ is good for modeling noble gases such as argon; electrically charged systems may add the Coulomb potential as well.
While pairwise potentials are conceptually simple, they are an important starting point for modeling atomic interactions. They are also often a constituent of more complex, multi-body potentials that are used to model more complicated systems that cannot be described by simple pairwise interactions alone. In a similar vein, the simplicity of pairwise potentials makes them an important starting place for porting efforts before moving on to conceptually and algorithmically complex potentials. We investigate the implementation of pairwise potentials in LAMMPS in greater detail in 4.1.
@@
The energy of the ReaxFF potential is a sum of traditional pairwise non-bonded interactions in which all neighboring atoms interact, empirical descriptions of two-, three-, and four-body bonded interactions in which only bonded sets of atoms interact, and terms encoding polarization and charge transfer effects. The charge transfer effects are encoded in a separate process known as the charge equilibration (QEq) method, defined by minimizing the total electrostatic energy of the system while maintaining charge conservation. Each portion of the energy is based on empirical observations of real dynamical systems. We investigate implementing and optimizing the critical portions of the ReaxFF potential in the Kokkos backend in 4.2.
The third case study is the machine-learning-based potential SNAP (Spectral Neighbor Analysis Potential) (Thompson et al., 2015).
Machine-learned potentials generally maintain the assumption of local interactions, but allow a more complex and general functional form that can consume information about the entire atomic neighborhood of an atom to produce the per-atom energy:
(2) $E=\sum_{i}E_{\text{ML}}\left(\left\{\vec{r}_{ik}\|r_{ik}<r_{\text{c}}\right\}\right).$
@@
We defer to (Thompson et al., 2015) for the more granular details of SNAP. We will introduce the minimum required mathematical formalism and then discuss the challenges and optimizations of the Kokkos implementation of SNAP in 4.3.

### v2

@@ Abstract.
Since its inception in 1995, LAMMPS has grown to be a world-class molecular dynamics code, with thousands of users, over one million lines of code, and multi-scale simulation capabilities. We discuss how LAMMPS has adapted to the modern heterogeneous computing landscape by integrating the Kokkos performance portability library into the existing C++ code. We investigate performance portability of simple pairwise, many-body reactive, and machine-learned force-field interatomic potentials. We present results on GPUs across different vendors and generations, and analyze performance trends, probing FLOPS throughput, memory bandwidths, cache capabilities, and thread-atomic operation performance. Finally, we demonstrate strong scaling on three exascale machines – OLCF Frontier, ALCF Aurora, and NNSA El Capitan – as well as on the CSCS Alps supercomputer, for the three potentials.
@@
copyright: ccconference: Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis (SC Workshops ’25), November 16–21, 2025, St Louis, MO, USAjournalyear: 2025doi: 10.1145/3731599.3767498isbn: 979-8-4007-1871-7/2025/11ccs: General and reference Performanceccs: Software and its engineering Ultra-large-scale systemsccs: Computing methodologies Parallel algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Distributed algorithmsccs: Computing methodologies Machine learningccs: Computing methodologies Molecular simulation
@@ 1. Introduction
Molecular dynamics is a method for studying materials and molecules by directly integrating Newton’s second law to obtain trajectories that describe chemical reactions, heat transport, diffusion, and other physical phenomena that occur at the time and length scales that are within computational reach of atomistic methods. Since the initial efforts by Rahman in 1964 (Rahman, 1964), the increase in applicability of molecular dynamics has been driven by the exponential growth in computational power that enables simulations of larger systems over longer timescales with higher accuracy (Plimpton and Thompson, 2012; Perez et al., 2025).
Further progress made in the 1990s and the spread of distributed computing presented a major leap. The natural spatial decomposition of molecular dynamics simulations made them a prime example for adopting the new massive parallelism paradigm, as demonstrated by the first releases of still-popular software such as LAMMPS in 1995 (Plimpton, 1995), as well as GROMACS (Abraham et al., 2015) and NAMD (Phillips et al., 2005). Over the next decade and a half, LAMMPS grew in scope and capability, from 23k lines of Fortran code in 1999, to 148k lines of C++ in 2009, but the MPI-distributed domain decomposition computing paradigm remained the same. Then, GPUs became more general purpose with the introduction of programming models such as CUDA. Their high FLOP rate and improved power efficiency gave significant advantages over CPUs, leading to widespread adoption of heterogeneous computing clusters.
With parallelism being a major strong point of GPUs, molecular dynamics codes with their large number of force calculations once again became an early adopter of the new heterogeneous computing paradigm. New codes were developed such as HOOMD-blue (Anderson et al., 2020) and HAL’s MD (Colberg and Höfling, 2011), while existing codes raced to adopt the new technology as quickly and efficiently as possible. For LAMMPS, whose modular package system (described in section 3.1 below) allows development of different approaches simultaneously, two competing strategies soon emerged.
@@
This paper instead focuses on the second strategy: the KOKKOS package in LAMMPS, which uses the Kokkos library (Trott et al., 2021) to achieve performance portability. The KOKKOS package (released in 2014) has design origins in the USER-CUDA package developed by Trott (Trott, 2011; Trott et al., 2010), which was released as part of LAMMPS in 2011. The USER-CUDA package attempted to be “GPU resident”, meaning as many kernels as possible run on the GPU. This avoids copying data between GPU and CPU at every timestep, leading to performance benefits. However, data transfer is still necessary to maintain compatibility when using functionality not yet ported to GPUs. The KOKKOS package brought the additional promise of keeping LAMMPS vendor-agnostic, which has proven vital given the modern introduction of large-scale computing resources using GPUs from AMD, Intel, and NVIDIA.
@@
In this paper, we discuss the current structure of LAMMPS and how the optional KOKKOS package provides seamless performance portability for the user when enabled. We include a comprehensive set of benchmarks for comparing how NVIDIA GPU performance has improved over hardware generations (V100-16GB-SXM3, A100-40GB-SXM4, and H100-HBM3-SXM5), and how LAMMPS benchmarks perform on three existing exascale systems: Frontier (AMD MI250X) (Top500, 2025d), El Capitan (AMD MI300A) (Top500, 2025b) and Aurora (Intel Data Center GPU Max 1550 – subsequently referred to as PVC) (Top500, 2024), along with the Alps supercomputer (NVIDIA Grace-Hopper Superchip – subsequently referred to as GH200) (Top500, 2025a) and NVIDIA’s Eos (NVIDIA DGX H100 Superpod) (Top500, 2025c). For reference, we include a comparison of different hardware in table 1.
Both AMD MI250X and Intel PVC contain two logical GPUs, and in this work single GPU results use a single “GCD” (GPU compute die) for MI250X and a single “stack” for PVC. We note that for single GPU results we focused on NVIDIA H100, while multi-GPU results used NVIDIA GH200. In the context of this work the differences are minimal; we quantify this in section appendix C.
We present case studies on the Lennard-Jones, SNAP, and ReaxFF potentials, and a discussion of how LAMMPS keeps up with current developments in the field of molecular dynamics such as machine-learning-based potentials.
@@
Figure 1. Mapping (red arrows) from user input script commands to LAMMPS class hierarchy and how the classes interact within and outside of the KOKKOS package. The pair styles operate on atomic data stored as pointers in AtomVecAtomic and Kokkos::DualViews in AtomVecAtomicKokkos, with the pointer in the former aliased to the CPU mirror of the latter. The EAM pair style requires additional communication, which is performed with calls to the LAMMPS communication classes. Following Unified Modeling Language (UML) conventions, solid black arrows indicate inheritance, while dashed arrows indicate usage.
@@
Table 1. GPU architecture properties, listing HBM memory bandwidth (“BW”) and capacity, floating point throughput (FP64, excludes matrix multiplication hardware), and cache size (single numbers are used for combined caches, two numbers for separate hardware and software managed caches). For AMD MI250X and Intel PVC a single logical GPU is listed, not the full package.
@@ 2. LAMMPS basic code structure
LAMMPS is one of the leading molecular dynamics simulation codes, in part because of the flexibility and collection of capabilities that have grown through over 30 years of development and user contributions (Kohlmeyer and Berger, 2025; Thompson et al., 2022).
@@ 2.2. Styles
These persistent classes are referred to as styles, which belong to different categories described more closely in the LAMMPS reference paper (Thompson et al., 2022). LAMMPS contains hundreds of different styles, in different categories. “Fix” styles have methods that are called at arbitrary points and intervals during the simulation to either modify the trajectory of the simulation or generate output, while “compute” styles can be viewed as restricted versions of fixes that only produce output accessible to the user within the input script, without modifying the state of the system. The case studies presented in section 4 are examples of pair styles, with each pair style implementing a different model for the potential energy, often referred to as simply “potentials”. At each timestep, the derivatives of the selected potential with respect to atomic coordinates provide the forces that are used for time integration, making pair styles the most important category of styles since they are typically the most expensive part of a simulation. The examples in section 4 span a range of complexities, from a simple pairwise potential to a state-of-the-art machine-learning-based potential.
3. LAMMPS-KOKKOS package
The KOKKOS package in LAMMPS was one of the first serious applications for Kokkos, with some of Kokkos’s design features directly inspired by use-cases from LAMMPS – including in the area of data structures and parallel execution. In this section, we describe how Kokkos is integrated to seamlessly enable performance portability without losing access to preexisting non-Kokkos functionality.
@@ 3.1. Packages and suffixes
The “accelerator” packages, on the other hand, provide replacement functionality in the form of accelerated styles, i.e., classes that provide equivalent functionality but improve computational performance on different hardware. This paper describes the KOKKOS package, which provides alternate versions of many classes that are accelerated by the Kokkos library for performance portability. As shown in figure 1, the KOKKOS package contains a PairEAMKokkos class for the EAM potential that is equivalent to the base PairEAM class from which it inherits, but its computational routines implement Kokkos library abstractions.
At runtime, the accelerated versions need to be selected through the map from input script command to C++ class described in section 2.1 and figure 1. Each accelerated style is registered and added to the map using the same macro as the non-accelerated style, with the convention that the accelerated styles append a package-specific suffix, namely /kk for the KOKKOS package. The user can then selectively choose the accelerated version of a style using, for example, the eam/kk pair style instead of eam in their input script, or they can use the Kokkos-accelerated version of all possible styles used in the input script by specifying a global suffix. Through careful design of the data structures and data movement patterns described in this section, the user is allowed to accelerate the parts of their simulation for which accelerated styles are available, without losing access to those for which no Kokkos support currently exists.
@@ 3.2. Data Structures
The primary data structure in Kokkos is the Kokkos::View, a structure that can represent multi-dimensional arrays, and encodes both data layout and data accessible through type mechanism via its Layout and MemorySpace template parameters. Kokkos::View strongly informed the design of the ISO C++23 std::mdspan capability. The Kokkos variants of styles in LAMMPS generally contain host and device variants of data encapsulated in a Kokkos::DualView. This data structure contains a compatible Kokkos::View for both a host (i.e. CPU) and a device (i.e. GPU) memory space to help with unstructured data synchronization. In particular, it has functionality to keep track of when data was modified, and thus when data has to be synced. The classic non-Kokkos data fields in LAMMPS are initialized to alias the allocations underlying the host View of the Kokkos data structures via Kokkos’s View interoperability with raw pointers.
@@
Another higher level data structure relied upon by LAMMPS is ScatterView. This data structure was designed to handle unstructured accumulation of data from multiple threads in a way that write conflicts are avoided. It can transparently swap between using atomic operations, a data duplication strategy, or even simple sequential accumulation in case of non-threaded execution. This is important to handle the different amounts of concurrency and atomic operation throughput on various architectures. On CPUs, data duplication with a subsequent combining step is often the most effective way to deal with write conflicts, while on GPUs data duplication is infeasible due to the large number of active threads (O(100,000)) and thus atomic operations need to be used.
@@ 3.3. Execution Control
Nearly every Kokkos-based style in LAMMPS is templated on the device type and is instantiated for both the default host and device execution space simultaneously. This dual instantiation pattern enables users to request a Kokkos-based style for either host CPU or GPU device at runtime from the input script through the /kk/host suffix and the /kk/device suffix (equivalent to /kk, see section 3.1), which can greatly improve execution control. One example of potential benefits is in the communication phase, where, depending on the size of the problem, the availability of GPU-aware MPI, and the actual hardware, it may be more performant to keep all communication routines (packing, unpacking, sending data) on host, or execute it on the device.
Kokkos parallel execution is controlled via execution patterns and execution policies. The primary patterns are parallel_for, parallel_reduce and parallel_scan which execute a callable for each index in an iteration space, and if applicable, combine results via reductions. The main policies are RangePolicy, MDRangePolicy and TeamPolicy. RangePolicy and MDRangePolicy simply iterate over a 1D or multi-dimensional range of indices, respectively. The latter also enables tiling, which can be beneficial to achieve better cache locality in multi-dimensional loop patterns.
TeamPolicy enables the use of hierarchical (nested) parallelism, i.e. exposing concurrency in non-tightly nested loops. Specifically it exposes a concept of thread teams, which can collaboratively execute nested parallel loops, as well as vector-level parallelism as a third nesting level using TeamThreadRange, TeamVectorRange, and ThreadVectorRange respectively. It also enables the use of scratch memory, which on GPUs can be mapped to software-managed caches such as NVIDIA’s “shared memory”, a topic which will be explored in more depth in section 4.4.
@@ 4. Case studies
For a general pairwise potential, the total energy of the system is a sum over all pairs of atoms that are within a specified cutoff distance. The interatomic forces are then the pairwise sum of the derivative of the potential energy with respect to position. The summation over pairs of close atoms is enabled by the neighbor list constructed and periodically updated by LAMMPS according to the algorithms described in the LAMMPS reference paper (Thompson et al., 2022).
@@
(1) $E=\sum_{i<k,r_{ik}<r_{\text{c}}}4\varepsilon\left[\left(\frac{\sigma}{r_{ik}}\right)^{12}-\left(\frac{\sigma}{r_{ik}}\right)^{6}\right]=\sum_{i<k,r_{ik}<r_{\text{c}}}U^{(2)}~(r_{ik}),$
with $r_{ik}$ the distance between atoms $i$ and $k$, $r_{\text{c}}$ is the interaction cutoff distance that may vary between atomic systems, and $\varepsilon$ and $\sigma$ are coefficients unique to LJ that may also vary between given atomic systems. LJ refers to a specific functional form of the more general idea of a two body potential $U^{(2)}(r_{ik})$. Different systems call for different functional forms. LJ is good for modeling noble gases such as argon; electrically charged systems may add the Coulomb potential as well.
While pairwise potentials are conceptually simple, they are an important starting point for modeling atomic interactions. They are also often a constituent of more complex, multi-body potentials that are used to model more complicated systems that cannot be described by simple pairwise interactions alone. In a similar vein, the simplicity of pairwise potentials makes them an important starting place for porting efforts before moving on to conceptually and algorithmically complex potentials. We investigate the implementation of pairwise potentials in LAMMPS in greater detail in section 4.1.
@@
The energy of the ReaxFF potential is a sum of traditional pairwise non-bonded interactions in which all neighboring atoms interact, empirical descriptions of two-, three-, and four-body bonded interactions in which only bonded sets of atoms interact, and terms encoding polarization and charge transfer effects. The charge transfer effects are encoded in a separate process known as the charge equilibration (QEq) method, defined by minimizing the total electrostatic energy of the system while maintaining charge conservation. Each portion of the energy is based on empirical observations of real dynamical systems. We investigate implementing and optimizing the critical portions of the ReaxFF potential in the KOKKOS package in section 4.2.
The third case study is the machine-learning-based potential SNAP (Spectral Neighbor Analysis Potential) (Thompson et al., 2015). Machine-learned potentials generally maintain the assumption of local interactions, but allow a more complex and general functional form that can consume information about the entire atomic neighborhood of an atom to produce the per-atom energy, $E_{i}=E_{\text{ML}}\left(\left\{\vec{r}_{ik}\|r_{ik}<r_{\text{c}}\right\}\right),$ which is then summed to the total energy.
@@
We defer to (Thompson et al., 2015) for the more granular details of SNAP. We will introduce the minimum required mathematical formalism and then discuss the challenges and optimizations of the Kokkos implementation of SNAP in section 4.3.